### v1

# Instruction-Level Weight Shaping: A Framework for Self-Improving AI Agents

Abstract
Large language models (LLMs) excel at surface fluency yet remain structurally static after pre-training; new or evolving domain knowledge is typically bolted on via retrieval-augmented generation (RAG) or parameter fine-tuning. In practice, RAG often retrieves facts without integrating them logically, adds latency and engineering overhead. Free-form prompt injection and ad hoc prompt engineering are brittle, prone to context-window drift, and can conflict with pre-trained knowledge. Fine-tuning, while effective for specific domains, is resource-intensive and risks catastrophic forgetting.
We propose Instruction-Level Weight Shaping (ILWS), which treats curated system instructions as external, auditable pseudo-parameters updated post-session via reflection and user feedback. After each session an LLM-driven Reflection Engine inspects the conversation trace, diagnoses reasoning successes or failures, and proposes typed deltas $\Delta K=(\Delta S,\Delta U,\Delta T)$ over instructions, user preferences, and tools. Each delta is version-controlled, evaluated under a sliding-window analysis of 1–5 star ratings, automatically repaired on first failure, and rolled back on repeated failure. When the accumulated edit budget crosses a threshold, the agent compiles a rating-weighted synthetic dataset and distils matured instruction-space gains into parameters, converting prompt-space improvements into weight-space without downtime.
[…]
1 Introduction
This paper introduces Instruction-Level Weight Shaping (ILWS), a lightweight framework for continual self-improvement in large language models (LLMs). In production systems, system instructions are treated as authoritative directives. ILWS reinterprets these instructions not as fixed configuration, but as a mutable, externalised memory channel, a low-cost, auditable surrogate for the model’s internal weights. Rather than updating parameters through costly fine-tuning or repeatedly fetching context via retrieval-augmented generation (RAG), ILWS uses a post-session reflection and feedback loop to produce knowledge deltas. These fine-grained edits gradually evolve the system prompt to better capture a domain’s logic, tools, and user expectations.
We contend that a large fraction of operational evolution can be handled by structured edits to the system prompt itself, provided edits are (i) feedback-driven and quantitative, (ii) reversible under governance, and (iii) recorded with code-like rigour. After each session, a stochastic Reflection Engine proposes a knowledge delta $\Delta K_{t}=(\Delta S_{t},\Delta U_{t},\Delta T_{t})$, which is trialled, score-gated (accepted only if a sliding-window rating improves by at least $\tau$ with significance $\alpha$), possibly repaired, and either accepted into a versioned knowledge state or rolled back. When the cumulative instruction budget exceeds a threshold, ILWS synthesises a rating-weighted dataset and distils persistent instruction-space edits into model weights (cf. Optional distillation, Eq. (3.4)).
Formally, at turn $t$ the agent uses a frozen backbone $f_{\theta}$ and a composite knowledge state $K_{t}=(S_{t},U_{t},T_{t})$:
$\hat{y}_{t}=f_{\theta}(x_{t},K_{t}),\qquad K_{t+1}=(S_{t}\oplus\Delta S_{t},\ U_{t}\oplus\Delta U_{t},\ T_{t}\oplus\Delta T_{t}).$ (1.1)
Here, $\oplus$ denotes applying the typed edit list to each component (insert/modify/delete). We show how this operational loop provides an explicit, auditable analogue of implicit low-rank weight shaping induced by prompts in transformer blocks (akin to low-rank adapters (LoRA) produced on-the-fly by prompts [5, 9]), and how it complements RAG and fine-tuning in practice.
Contributions.
  • Formalise ILWS with typed deltas and governance. We cast structured prompt edits over $K=(S,U,T)$ as explicit, version-controlled surrogates for low-rank weight updates, with git-backed persistence and rollback.
  • Statistical gate with repair/rollback. We introduce a sliding-window, statistically grounded acceptance rule parameterised by $(\tau,\alpha)$, admitting edits only if average ratings improve by at least $\tau$ with significance $\alpha$, with one-shot repair and rollback on second failure.
  • Theory link to implicit low-rank shaping. We articulate how instruction edits influence prompt-conditioned activations to yield LoRA/IA3-like low-rank effects, providing an explicit, auditable analogue of implicit updates induced by prompts.
  • Production architecture with autonomous tool synthesis. We present a stochastic reflection loop, sandboxed tool generation and integration, and operational guardrails (policy invariants, statistical change detection on ratings).
[…]
2 Theoretical context and related work
Implicit low-rank updates from context.
A transformer block comprising a contextual layer followed by an MLP can be viewed as applying an implicit rank-one update to the first MLP layer, computed from the context tokens; iteratively consuming tokens yields an online gradient-like dynamic over an effective weight matrix [2]. ILWS makes such influences explicit by editing instructions outside the network, preserving auditability and persistence.
Path-kernel view of trained models.
Models trained by gradient descent are approximately kernel machines whose predictions can be expressed via path kernels that integrate gradient similarities along the optimization trajectory [3]. Under this view, retrieved or prompted examples act as non-parametric supports. Selecting context that aligns with gradient-path features should improve reliability—a perspective we exploit in our retrieval-free but instruction-grounded edits.
RAG and modular augmentation.
RAG is effective but requires careful indexing, retrieval granularity, reranking, compression, and adaptive triggering to avoid irrelevant or counterfactual context [4]. ILWS sidesteps per-call retrieval by baking vetted rules and patterns into instructions, while remaining compatible with optional RAG when needed.
Self-improvement and reflection.
Memory-augmented prompting (e.g., MemPrompt [10]), reflection-based retries (e.g., Reflexion [12]), and active retrieval (e.g., Self-RAG, FLARE [1, 6]) demonstrate gains from critique and adaptation. ILWS unifies these ideas post-session, gating durable edits with human-centric scores and governance, and adding autonomous tool synthesis.
LLM tool-use and modular retrieval.
Orthogonal lines of work teach models to call tools (e.g., Toolformer) and to interleave retrieval and generation (e.g., Demonstrate-Search-Predict, active retrieval) [11, 7, 6, 1]. These systems improve evidence acquisition at inference time, but typically do not persist successful reasoning as durable edits. ILWS is complementary: it converts recurrent patterns into authoritative instructions and (optionally) new tools, reducing online retrieval while preserving auditability.
3 Problem formulation
Let $f_{\theta}$ be a frozen LLM backbone. The agent maintains a knowledge state
$K_{t}=(S_{t},U_{t},T_{t}),$ (3.1)
where $S_{t}$ is the current instruction set, $U_{t}$ captures user learnings or preferences, and $T_{t}$ is the registry of callable tools. During inference the system prompt is constructed deterministically from $K_{t}$ and provided to $f_{\theta}$ together with the user input $x_{t}$; cf. Eq. (1.1). At session end the Reflection Engine emits a candidate $\Delta K_{t}=(\Delta S_{t},\Delta U_{t},\Delta T_{t})$.
Score-gated acceptance.
Let $r_{i}\in\{1,\ldots,5\}$ be user ratings (5-point Likert). A candidate edit $\Delta K_{t}$ is provisionally deployed immediately and evaluated over the next $N_{\text{win}}$ interactions. Define sliding-window means over $N_{\text{win}}$ interactions:
$\bar{r}_{\mathrm{prev}}=\frac{1}{N_{\text{win}}}\sum\nolimits_{i=t-N_{\text{win}}}^{t-1}r_{i},\qquad\bar{r}_{\mathrm{new}}=\frac{1}{N_{\text{win}}}\sum\nolimits_{i=t}^{t+N_{\text{win}}-1}r_{i}.$ (3.2)
We accept the provisional edit if
$\bar{r}_{\mathrm{new}}\geq\bar{r}_{\mathrm{prev}}+\tau\quad\text{and}\quad p\text{-value}\leq\alpha,$ (3.3)
where the $p$-value comes from a one-sided Welch $t$-test by default (falling back to Mann–Whitney if normality fails Shapiro–Wilk), and typical hyper-parameters are $(\tau,\alpha)=(0.05,0.05)$. Note that $\tau=0.05$ corresponds to one-twentieth of a star on a 5-point scale. Equivalently, the gate admits an edit only if it improves the average rating by at least $\tau$ with significance $\alpha$. On first failure we solicit a typed automatic repair $\Delta K^{\prime}_{t}=(\Delta S^{\prime}_{t},\Delta U^{\prime}_{t},\Delta T^{\prime}_{t})$; on a second consecutive failure we roll back to the last tagged good state.
Optional distillation.
We track a running instruction-change budget
$B_{T}=\sum_{i=1}^{T}\big{(}\|\Delta S_{i}\|+\|\Delta U_{i}\|+\|\Delta T_{i}\|\big{)},$
where $\|\cdot\|$ is a simple size metric such as token count or edit length. When $B_{T}\geq M$, we synthesise a rating-weighted dataset $D_{\text{syn}}=\{(x,K,\hat{y},r)\}$ and solve
$\theta^{\star}\in\arg\min_{\theta^{\prime}}\sum_{(x,K,y)\in D_{\text{syn}}}w(x,y)\,\mathcal{L}_{\mathrm{CE}}\!\left(f_{\theta^{\prime}}(x,K),y\right)\quad\text{(token-level cross-entropy)},\qquad w(x,y)=\frac{r-1}{4}\in[0,1],$ (3.4)
then redeploy $f_{\theta^{\star}}$ and reset $B_{T}$.
4 Instruction-Level Weight Shaping (ILWS)
Figure 1 (left-to-right) depicts the four phases.
4.1 Phase 1: Inference
Given $(x_{t},K_{t})$, the agent returns $\hat{y}_{t}=f_{\theta}(x_{t},K_{t})$ per Eq. (1.1). In our reference system, $K_{t}$’s components are versioned, serialised JSON fragments deterministically composed into the system prompt; the schema is version-pinned so edits remain diff-friendly. The session stores a transcript, tool logs, and the rating $r_{t}$.
4.2 Phase 2: Post-session reflection and update
A session-end hook or cron job invokes the Reflection Engine $R$ with transcript, tool logs, and recent ratings. $R$ emits a typed delta
$\Delta K_{t}=(\Delta S_{t},\Delta U_{t},\Delta T_{t})$
expressed as calls into a self-modification API (appendInstruction, modifyInstruction, createTool, deprecateTool, addUserPreference, …) plus a structured rationale (YAML diagnostics, score deltas). The candidate is applied immediately and evaluated over the next $N_{\text{win}}$ sessions under the score-gating rule (3.3).
Autonomous tool synthesis.
If $\Delta T_{t}\neq\varnothing$, the Tool Manager compiles and unit-tests generated Python in a sandbox (networkless, no egress; OCI/seccomp profile). On success, the tool signature is appended to $T$, and a concise usage rubric is inserted into $S$ to make the capability discoverable by the model; admins may revert within a review window $\Delta t$.
4.3 Phase 3: Persistence and governance
Accepted deltas are committed to an immutable git repository and tagged with a knowledge checkpoint; admins act as observers and may optionally veto by triggering a one-click revert within a configurable review window $\Delta t$. A dashboard exposes diffs, sliding-window metrics, gate decision traces $(\tau,\alpha,p\text{-value})$, confidence intervals, failure analyses, and one-click reverts. Human veto flags are recorded and fed back into subsequent reflection prompts to prevent repeated proposals that were explicitly declined.
4.4 Phase 4: Long-term evolution
When the instruction token budget exceeds $M$, ILWS synthesises $D_{\text{syn}}$ and distils to weights via Eq. (3.4). Fine-tuning runs offline; live traffic continues to hit the frozen $f_{\theta}$. This collapses stable prompt-space edits into parameters, freeing context for future growth and keeping the JSON prompt under a budget $C$ tokens to avoid context-window bloat.
Figure 1: ILWS data flow with four phases: inference, self-improvement, persistence/governance, and long-term evolution (distillation). The right panel expands the reflection/score-gating/repair/rollback logic.
4.5 Algorithms
Algorithm 1 Session loop with ILWS
[…]
2:Initialize $K_{0}=(S_{0},U_{0},T_{0})$, instruction-change budget $B\leftarrow 0$ $\triangleright$ edit-budget $\|\Delta S\|{+}\|\Delta U\|{+}\|\Delta T\|$
3:Initialize rating buffer; $\bar{r}_{\mathrm{prev}}\leftarrow 3$ $\triangleright$ neutral prior
4:Set window size $N_{\text{win}}$ and margins $(\tau,\alpha)$.
5:for session $t=1,2,\ldots$ do
6:  Receive input $x_{t}$; produce $\hat{y}_{t}=f_{\theta}(x_{t},K_{t})$; log transcript and tools; obtain rating $r_{t}$; append $r_{t}$ to sliding-window rating buffer (last $N_{\text{win}}$).
7:  if buffer length $N_{\text{win}}$ then
8:   $K_{t+1}\leftarrow K_{t}$ $\triangleright$ state unchanged during warm-up
9:   continue $\triangleright$ warm-up: skip gating until buffer is full
10:  end if
11:  $\Delta K_{t}\leftarrow R(\text{transcript},\text{tools},\text{buffer})$ $\triangleright$ Reflection Engine uses $r_{t-N_{\text{win}}:t-1}$
12:  Provisionally set $K_{t}^{\mathrm{tmp}}\leftarrow(S_{t}\oplus\Delta S_{t},\ U_{t}\oplus\Delta U_{t},\ T_{t}\oplus\Delta T_{t})$ $\triangleright$ evaluate asynchronously (scheduled task; does not block inference) over next $N_{\text{win}}$ sessions
13:  $r_{\text{prev}}\leftarrow r_{t-N_{\text{win}}:t-1}$; $r_{\text{new}}\leftarrow r_{t:t+N_{\text{win}}-1}$ $\triangleright$ rating windows
14:  $\bar{r}_{\mathrm{prev}}\leftarrow\mathrm{mean}(r_{\text{prev}})$; $\bar{r}_{\mathrm{new}}\leftarrow\mathrm{mean}(r_{\text{new}})$
15:
16:  $p\leftarrow\textsc{WelchTTest}(r_{\text{prev}},\,r_{\text{new}})$ $\triangleright$ one-sided; fallback Mann–Whitney; runs after window closes
17:  if $\bar{r}_{\mathrm{new}}\geq\bar{r}_{\mathrm{prev}}+\tau$ and $p\leq\alpha$ then
18:   $K_{t+1}\leftarrow K_{t}^{\mathrm{tmp}}$; commit & tag in git; $B\leftarrow B+\|\Delta S_{t}\|+\|\Delta U_{t}\|+\|\Delta T_{t}\|$.
19:   $\triangleright$ Admin may veto within review window $\Delta t$
20:   if veto then
21:     rollback to last tag; $K_{t+1}\leftarrow K_{t}$; $B\leftarrow B-\bigl{(}\|\Delta S_{t}\|+\|\Delta U_{t}\|+\|\Delta T_{t}\|\bigr{)}$ $\triangleright$ revert commit if veto arrives later
22:   else
23:     $\triangleright$ no veto, budget check
24:     if $B\geq M$ then
25:      Distill via Eq. (3.4); deploy $f_{\theta^{\star}}$; $B\leftarrow 0$
26:     end if
27:   end if
28:  else
29:   Request typed repair $\Delta K_{t}^{\prime}=(\Delta S_{t}^{\prime},\Delta U_{t}^{\prime},\Delta T_{t}^{\prime})$ from $R$; re-evaluate once under the same acceptance gate. $\triangleright$ if repair is accepted, follow the success path above
30:   If still failing, rollback to last tag; set $K_{t+1}\leftarrow K_{t}$; $\triangleright$ discard failed candidate
31:  end if
32:end for
4.6 Safety guardrails
We separate implemented guardrails in the reference PoC from recommended hardening for production. The gate in Eq. (3.3) and the repair/rollback loop already constrain behavioural drift; below we focus on code and data safety.
Implemented in the PoC.
  • Version control and audit. All edits to instructions and tools are committed to git with timestamps; JSON saves create on-disk backups; reflection prompts and outcomes are written to an audit log directory.
  • Knowledge tagging and rollback. The feedback module can tag the current knowledge state and revert to the last good tag under degradation, keeping a backup of the faulty state for review.
  • Tool creation denylist. Generated tool code is scanned for a deny-list of dangerous strings (e.g., sudo, chmod, curl, wget, eval() both in code and in file-like parameters before execution.
  • Path isolation. Each tool executes with its working directory switched to a per-tool sandbox folder; file helpers storeInDisk/loadFromDisk validate filenames to prevent path traversal and absolute paths.
Recommended for production.
  • Static analysis, not regex alone. Parse generated code with an AST and enforce an allow-list of modules (e.g., math, json); block imports such as os, subprocess, socket, requests, dynamic eval/exec, and file I/O outside provided helpers.
  • Runtime isolation. Run tools in a networkless, seccomp-filtered OCI container (with an AppArmor profile) with resource limits (CPU, memory), execution timeouts, and capped output size.
  • Unit-test gate. Require a minimal unit-test scaffold to pass before registering a tool; on failure, quarantine the tool and surface the failure diff in the dashboard.
  • Secret and egress checks. Scan code and outputs for credential patterns and block environment access; maintain an explicit policy for any outbound calls (default: none).
  • Explicit approval rubrics for side-effects. For tools that mutate external state, insert an approval rubric into $S$ (what to confirm, with which fields) and require an explicit, structured user confirmation step.
  • Edit scope and allowlists. Constrain $\Delta S$ to edit only within pre-declared sections (global/product/tenant) and enforce allowlisted patterns for sensitive policies.
  • Statistical change detection (optional). Lightweight EWMA/CUSUM on ratings can flag sudden drops even if the gate passes; this remains a secondary signal to user ratings.
4.7 Framework versus reference implementation
ILWS is a framework: it governs durable behavioural knowledge via a typed state $K=(S,U,T)$ and audited edits $\Delta K=(\Delta S,\Delta U,\Delta T)$. Edits are proposed post-session by a reflection process, screened by a statistically grounded gate, repaired once on failure, rolled back on repeated failure, and periodically distilled when an edit budget is exceeded. These steps are conceptual and agnostic to any specific rating signal, statistical test, or sandbox.
To avoid conflating the framework with one instantiation, we adopt a minimal validity contract. A substitute mechanism remains compliant if (i) the feedback signal (human or automated) is monotonically correlated with task quality, (ii) the gate compares pre- and post-edit quality on at least $N_{\min}$ samples and controls family-wise Type I error at $\leq\alpha$, and (iii) the execution sandbox enforces policy-governed network egress (default-deny with explicit allowlists and audit logging), restricts file I/O to a declared workspace, and blocks dynamic code execution unless explicitly whitelisted. Any mechanism satisfying these conditions can replace the defaults without altering ILWS.
Our reference system meets this contract with pragmatic defaults: five-star user ratings; a sliding-window one-sided Welch $t$-test gate with $(\tau,\alpha)$ and window size $N_{\text{win}}$; a one-shot typed repair followed by rollback; a deny-list sandbox; optional autonomous tool synthesis; and an edit-budget trigger for offline distillation. These are defaults, not intrinsic to ILWS, and can be swapped for valid alternatives such as reward-model scores, Bayesian or sequential tests, AST/OPA policy checks, or a seccomp-filtered OCI container.
Deployments may also inject ephemeral context at inference time (for example, server-resources, transaction analyses, or per-ticket metadata). In the “L0 Support” prototype, such telemetry aids diagnosis but is not persisted in $K$. Only rubrics, preferences, or tools that survive the gate become durable edits, preserving auditability and simplifying multi-tenancy.
5 Why instruction edits behave like weight shaping
Let $T_{W}=M_{W}\circ A$ denote a contextual layer $A$ followed by an MLP with first-layer weight $W$ and the remaining network $f_{\theta}$. For a token $x$ with context $C$ (which includes instruction tokens $S$), Dherin et al. [2] show—empirically and via linearised analysis—that the output with $C$ can be approximated by the output without $C$ but with a rank-one perturbation $\Delta W(C)$ applied to the first MLP layer (see their derivations/experiments). Hence the map $C\mapsto\Delta W(C)$ factors through the contextual representation $a(C):=A(x;C)$.
Consider a small edit $\delta S$ to the instruction tokens inside $C$. If $a(C)$ is $L_{S}$-Lipschitz in $S$ locally (measured in the $\ell_{2}$ norm of token-embedding differences), and the map $\Phi:a\mapsto\Delta W$ is smooth with local operator-norm bound $\kappa:=\|\nabla_{a}\Phi(a(C))\|_{\mathrm{op}}$ near $a(C)$, then
$\left\|\Delta W\big{(}C[\,S\leftarrow S{+}\delta S\,]\big{)}-\Delta W(C)\right\|_{F}$ $=\left\|\Phi\!\left(a\!\left(C[\,S\leftarrow S{+}\delta S\,]\right)\right)-\Phi\!\big{(}a(C)\big{)}\right\|_{F}$ (5.1)
$\leq\kappa\,\left\|a\!\left(C[\,S\leftarrow S{+}\delta S\,]\right)-a(C)\right\|_{2}$
$\leq\kappa L_{S}\,\|\delta S\|_{2}.$
Thus, under these local smoothness assumptions, an edit $\delta S$ scales the induced update by at most $\kappa L_{S}\,\|\delta S\|_{2}$. Small, structured edits to $S$ therefore act as a controlled dial on the magnitude (via $\kappa L_{S}$) and direction (through $a(C)$) of the effective low-rank update. This argument is local and qualitative: it relies on the rank-one approximation and smoothness in a neighbourhood of the current context and does not claim weight-level equivalence to fine-tuning. Because $\kappa L_{S}$ upper-bounds the update magnitude, ILWS’s $(\tau,\alpha)$ gate implicitly limits effective weight drift. Large edits and cross-token interactions beyond this neighbourhood fall outside the bound and are empirically screened by the gate.
We note that standard dot-product attention is not globally Lipschitz [8]; hence no uniform bound can be guaranteed. Our argument is therefore local, and ILWS enforces stability operationally via its $(\tau,\alpha)$ score-gate with repair/rollback, which screens out edits that would induce disproportionate drift.
The path-kernel view [3] complements this picture: path kernels interpret a trained network as a linear model in an implicit kernel defined by gradient paths. Edits that steer $a(C)$ toward features aligned with those paths induce low-rank tweaks that better match the desired behaviour; ILWS operationalises this by proposing instruction edits and accepting them only under the score-gated objective.
6 Experimental Evaluation
This section details a longitudinal, single-operator study of the ILWS framework embodied in a proof-of-concept tool named "L0 Support." The study was conducted within a live Adobe Commerce Cloud Level 2/Tier-3 support engineering environment. All performance data reflects the real-world ticket resolution throughput of the author.
6.1 Experimental Setting and Baselines
The primary evaluation environment is high-stakes technical support, where correctness, precision, and efficiency are paramount. The operator’s role involves diagnosing and resolving complex performance and configuration issues on the Adobe Commerce platform.
We establish two key baselines:
  1. i)
    Manual Throughput: The operator’s historical performance without any AI assistance. The established average was 50 resolved tickets per month, working a standard full-time schedule. A high-effort attempt to clear a backlog, yielded a maximum of 90 tickets in one month, demonstrating a practical ceiling for manual work.
  2. ii)
    RAG: An initial approach using Retrieval-Augmented Generation was tested and discarded. While RAG could retrieve documentation and configurations, it failed to perform causal reasoning applicable to novel issues. It often regurgitated irrelevant data and introduced significant per-message latency, measured between 800ms and 2000ms, making it unsuitable for an interactive diagnostic workflow.
6.2 Implementation Details
The "L0 Support" PoC was developed to handle the large and evolving knowledge base of Adobe Commerce. Key implementation choices include:
  • Model Selection: While several models were evaluated, Google’s Gemini-2.5-pro demonstrated superior reasoning capabilities during the self-reflection phase. However, as fine-tuning is not yet available for this model, GPT-4.1 was selected in a second PoC to accommodate the distillation phase (Phase 4).
  • Context Window: The continuous accumulation of domain-specific instructions necessitated a model with a 1 million token context window to hold the evolving system prompt.
6.3 Quantitative Results
The introduction of the ILWS-powered tool resulted in a dramatic and sustained increase in operator throughput and efficiency, even with reduced working hours. All calculations are based on an average of 22 working days per month and a 7.5-hour standard workday.
In the first month of deployment, the operator resolved 120 tickets while working only part-time (mornings). This represents a 140% increase in total ticket volume compared to the 50-ticket full-time baseline, achieved in roughly half the working hours. A subsequent three-week (15-day) sprint saw the resolution of 100 tickets working only afternoons.
Throughput Analysis.
To normalize these results, we analyze performance in terms of Tickets Per Hour (TPH).
  • Baseline TPH: 50 tickets / (7.5 hours/day $\times$ 22 days) $\approx$ 0.30 TPH.
  • ILWS TPH: Over a recent two-day period, the operator consistently resolved 13 tickets per day while working an average of 6 hours. This yields a measured throughput of 13 tickets / 6 hours $\approx$ 2.17 TPH.
This represents a throughput increase of over 7.2 $\times$ (2.17 / 0.30) compared to the manual baseline. The productivity gains are summarized in Table 1.
Table 1: Performance Metrics
Metric Manual Baseline ILWS Performance Improvement Factor
[…]
Hours / Ticket $\sim$3.30 $\sim$0.46 $\sim$86% Reduction
[…]
6.4 Qualitative Analysis
Beyond raw throughput, ILWS demonstrated significant qualitative improvements in the diagnostic process.
Drastic Reduction in Hallucinations.
Initially, without the domain-specialized instructions, the model’s suggestions were helpful in only about 2 out of 10 zero-shots attempts, requiring multiple iterations of prompting and correction. After the ILWS system matured its instruction set, its suggestions for performance-related tickets were accurate enough to solve the issue in zero or one shot in 9 out of 10 cases. This marks a shift from a 20% to a 90% first-shot resolution rate for this ticket category, indicating a significant reduction in model hallucination and an increase in reasoning precision.
A Case Study.
A compelling example of the reflection mechanism occurred during a performance investigation. The model initially hypothesized that high memory consumption in ‘php-fpm‘ workers was caused by cron jobs. The operator provided a crucial correction:
"Cron jobs run on ‘php-cli‘, not ‘php-fpm‘. ‘php-fpm‘ serves web traffic from users, APIs, or bots."
The post-session reflection engine processed this feedback and proposed an update to its system instructions, adding the rule that ‘php-fpm‘ is exclusively for web traffic while ‘php-cli‘ handles background tasks like crons. When a new session was started to simulate the same issue from scratch, the model, now equipped with this new instruction, immediately and correctly identified the root cause as high request volume in its first response.
System Maturation.
Over the course of approximately 300 support sessions, the system underwent 80 distinct instruction updates. In the initial phases, the operator rolled back only around 25% of the modifications suggested by the reflection engine, demonstrating a high degree of relevance and accuracy in the system’s self-improvement proposals. This iterative process of refinement is directly responsible for the observed increase in precision and speed.
7 Discussion: ILWS vs. RAG vs. fine-tuning
Optimization view (brief).
ILWS can be seen as a gradient-like supervisory optimisation in instruction space: small edits $\Delta S$ are proposed and accepted only if a sliding-window rating objective improves (a trust-region-like gate), with repair/rollback acting as a line search and early stopping. This bandit/RL perspective complements the implicit low-rank view: instructions serve as persistent pseudo-weights shaping behaviour, while true backpropagation applies only during the optional distillation stage. RAG is effective for fast-changing, citable knowledge but adds latency and risks irrelevant context [4]. Fine-tuning imprints stable competencies but is costly to run continuously. ILWS covers the operational middle: authoritative instructions for last-mile rules, validated online and distilled offline. The link to implicit low-rank shaping suggests edits as controlled, low-rank function tweaks rather than sprawling heuristics.
Table 2: Positioning summary.
Dimension ILWS RAG FT
Online latency low medium low
Update cadence per session per call offline
Auditability high medium medium
Drift control score-gated retriever-tuned data/process
Cost low medium high
8 Limitations and risks
ILWS assumes access to ratings that correlate with quality; noisy or gamed feedback can misguide edits. Tool synthesis, if under-specified, can create unsafe actions; our sandbox and policy invariants reduce but do not eliminate this risk. Finally, the theory-to-practice link is qualitative: while instruction edits influence effective low-rank updates, quantifying alignment remains an open problem.
9 Conclusion
Instruction-Level Weight Shaping offers a lightweight, auditable path to continual improvement: treat system instructions as dynamic surrogates for weight updates, gated by human feedback, governed like code, and periodically distilled. Empirically, ILWS delivered multi-fold productivity gains and fewer hallucinations without online retrieval or constant fine-tuning. The framework operationalizes emerging theory on in-context weight shaping and connects it to practical agent engineering.